Groundwater is softened to 80 mg/L hardness as CaC
Answer #1

LIME ADDITION Hardness Lime Precipitate CO2 Ca(HCO3)2Ca(OH)2->2CaCO32H20 Mg(HCO3)2Ca(OH ->CaCO+ MgCO 2H20 MgCO CaO >CaCO+ Mg(

Lime dosage = The carbon dioxide as CaCO3 + Ca2+ concentration

= 2 + 100 = 102 mg/L    final hardness is 80 therefore the dosage is 102 - 80 = 22 mg/L CaCO3

Soda ash dose = (Non carbonate hardness as CaCO3 * Calcium as carbonate) / Mg = 260*100/8 = 3250 mg/L